When Monica first entered the workforce in 2000 her annual salary was $41,500. Since that time Monica has been promoted. Now Monica wants to compare her income progression in real terms with her current salary in of $115,000. If CPI was 136 in 2000 and 156 in 2016, what is the difference between her starting salary and her current salary in 2016 price equivalents?
1
Expert's answer
2018-03-23T09:06:07-0400
2016CPI/2000CPI = 156/136 = 1.15
This ratio tells us that one dollar in 2000 is equivalent to $1.15 in 2016.
Therefore, the real Monica’s salary in 2016 is: $115,000/1.15 = $100,000
The difference is: $100,000-$41,500 = $58,500